in two years time, the square of Bill age then will equal four times the square of is age now. How old will he be in three years time?
Answer by lwsshak3(11628)

in two years time, the square of Bill age then will equal four times the square of is age now. How old will he be in three years time?
***
let x=Bill's present age
In 2-yrs time:
x+2=Bill's age
..
(x+2)^2=4x^2
x^2+4x+4=4x^2
3x^2-4x-4=0
(3x+2)(x-2)=0
x=-2/3 (reject)
or
x=2(Bill's present age)
In 3 yrs:
x+3=5
How old will Bill be in three years time? 5